OpenStudy (anonymous):

Help with 3 - 2x = -5 possible answers: 2 4 -2 1/5 I see none of these as possible answers as none of them equal -5. Am I doing something wrong, or is the equation inaccurate?

OpenStudy (photon336):

let's see

OpenStudy (photon336):

3-2x = -5

OpenStudy (photon336):

the first thing we need to do is isolate x we do this by subtracting 3 Do you see why I did this? 2x = 3-(-5)

OpenStudy (nethercreep333):

Let's solve your equation step-by-step. 3−2x=−5 Step 1: Simplify both sides of the equation. 3−2x=−5 Simplify: (Show steps) −2x+3=−5 Step 2: Subtract 3 from both sides. −2x+3−3=−5−3 −2x=−8 Step 3: Divide both sides by -2. −2x −2 = −8 −2 x=4

OpenStudy (photon336):

@NetherCreep333 I did it a differently 2x = 3-(-5) 2x = 8 x =4

OpenStudy (anonymous):

Answer would be 4, 3-(-5) = 8, 2x4 = 8. Makes sense I guess.

OpenStudy (photon336):

@BillGatessDanceMoves try one on your own like this 4-4x = -12

OpenStudy (photon336):

solve this for me based on what I did

OpenStudy (anonymous):

Well the way I usually do it is just take the possible answers and do the math until I find the correct one. I did it like this to find the answer for that. 4x(-2) = -8, -8 -4 = -12

OpenStudy (photon336):

then you will understand how to do it yourself

OpenStudy (photon336):

wait so what's your answer?

OpenStudy (anonymous):

-2.

OpenStudy (photon336):

@BillGatessDanceMoves let's check your answer

OpenStudy (photon336):